X^2 ydx + (y +x^3)dy=0 Find the general solution of differential equations.
1
Expert's answer
2020-08-23T17:19:16-0400
"Solution"

"x^2 ydx + (y +x^3)dy=0 \implies x^2ydx=-(y +x^3)dy\\
\frac{dx}{dy}=\frac{-y-x^3}{x^2y}\\
\frac{dx}{dy}=-\frac{1}{x^2}-\frac{x}{y}\implies \frac{dx}{dy}+ \frac{x}{y}= -\frac{1}{x^2}\\
\frac{dx}{dy}+ Rx= S\\
R=\frac{1}{y},\ S=-\frac{1}{x^2}\\
IF=\epsilon^{\int Rdy}=\epsilon^{\int \frac{1}{y}dy}=\epsilon^{log\ y}=y\\
x \cdot IF = \int S \cdot IF\ dy+C\\
x \cdot y = \int -\frac{1}{x^2} \cdot y\ dy+C= \int -\frac{y}{x^2}\ dy+C= -\frac{y^2}{2x^2}+C\\
\therefore xy=-\frac{y^2}{2x^2}+C-------->Answer"